More Hits Company manufactures aluminum baseball bats that it sells to university athletic departments. It has developed the following per unit standard costs for 2019 for each baseball bat: Direct Materials Direct Labor Manufacturing Overhead Standard quantity 2 pounds (Aluminum) ½ hour ½ hour Standard Price $4.00 $10.00 $6.00 Unit standard cost $8.00 $5.00 $3.00 In 2019, the company planned to produce 120,000 baseball bats at a level of 60,000 hours of direct labor. Actual results for 2019 are presented below: 1. Direct materials purchases were 246,000 pounds of aluminum which cost $1,020,900 2. Direct materials used were 220,000 pounds of aluminum. 3. Direct labor costs were $575,260 for 58,700 direct labor hours actually worked. 4. Total manufacturing overhead was $352,000 5. Actual production was 114,000 baseball bats Instructions Compute the following variances: 1. Direct materials price. 2. Direct materials quantity. 3. Direct labor price. 4. Direct Labor quantity. 5. Total overhead variance.
Answer #1
First let's arrange the data properly.
Direct Material Direct Labor Manufacturing OH
Standard Quantity 2 pounds 1/2 hour 1/2 hour
Standard Price $4.00 $10.00 $6.00
Unit Standard Cost $8.00 $5.00 $3.00
1 Material Price variance = (Actual Price - Standard price)*Actual quantity
Material Price variance = ($4.15 - $4.00)246,000 pounds =$36,900(U)
2 Material quantity variance = Standard price(Actual Quantity - Standard Quantity)
Material quantity variance = $4.00(220,000 - 228,000) =$32,000(F)
Standard Quantity =114,000 bats * 2 pound =228,000 pound of aluminium
3 Labor Price Variance =Actual hours(Actual rate - Standard rate)
Labor Price Variance =58,700 hours($9.80 - $10.00) =$11,740(F)
4 Labor quantity variance =Standard Rate(Actual Hours - Standard Hours)
Labor quantity variance =$10.00(58,700 - 57,000 hours) =$17,000(U)
Standard Hours =114,000 bats*1/2 hours =57,000 hours
5 Total Overhead Variance =Actual Overhead - Overhead applied
Total Overhead Variance =$352,000 - $342,000 =$10,000(U)
Overhead applied =Standard hours*Standard price =57,000 hours * $6.00 =$342,000